For a standard normal distribution find the following

P(z > 1.35)

P(z < 2.28)

P(-1.01 < z < 1.3)

Find C for the following

P(z < c) = 0.4834

P(z > c) = 0.1793
Answer #1

P(z > 1.35) = 1 - P(z < 1.35)

P(z > 1.35) = 1 - 0.9115 ( from table )

P(z > 1.35) = 0.0885

P(z < 2.28) = 0.9887 ( from table )

P(-1.01 < z < 1.3) = P( z < 1.3 ) - P( z < -1.01 )

P(-1.01 < z < 1.3) = 0.9032 - 0.1562 ( from table )

P(-1.01 < z < 1.3) = 0.747

Find C for the following

P(z < c) = 0.4834

c = 0.04 ( from table )

P(z > c) = 0.1793

P(z < c) = 1 - 0.1793

P(z < c) = 0.8207

c = 0.92 ( from table )